OpenStudy (anonymous):

What is the area of a square with side lengths of 7 feet? 28 ft2 A. 28ft B. 28 ft2 C.49ft D.49ft2 MEDAL AND FAN PLEASE @amistre64 @j-bird

OpenStudy (anonymous):

Someone help please @amistre64 @phi @JFraser

OpenStudy (igreen):

The area of a square is \(s^2\), where s = side length. The side length here is 7 feet. So what's \(7^2\)?

OpenStudy (amistre64):

in this case we want the dimension squared as well

OpenStudy (anonymous):

49 RIGHT

OpenStudy (amistre64):

(7*feet)^2 = 7^2 * feet^2

OpenStudy (igreen):

Yep. @amistre64 just said we square the dimension too. So feet becomes feet^2

OpenStudy (anonymous):

so 49ft2 right

OpenStudy (anonymous):

@iGreen @amistre64 @amistre64 @Preetha @WolfMaster @Hero

OpenStudy (anonymous):

I think it is 49ft2

OpenStudy (anonymous):

49ft2

OpenStudy (igreen):

Yes, it's 49ft^2. We already said it is 49, and we already said it's ft^2.